# ID-Composer: Multi-Subject Video Synthesis with Hierarchical Identity Preservation

### v2

@@ 3.3 Training with Online Reinforcement Learning
$\mathcal{J}_{GRPO}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (3)
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{t}^{i}\left(\theta\right)\hat{A}_{i,t},\ \text{clip}\left(r_{t}^{i}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\,\|\,\pi_{\text{ref}}\big]\right\},$
$\text{where}\quad r_{t}^{i}\left(\theta\right)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
@@
the relative quality $\hat{A}_{i,t}$ of the i-th response is computed as
$\hat{A}_{i,t}=\frac{r_{i}-\text{Mean}(\{r_{1},r_{2},\ldots,r_{n}\})}{\text{Std}(\{r_{1},r_{2},\ldots,r_{n}\})},$
@@
Flow-GRPO extends GRPO to the rectified flow setting, while convert flow matching’s determinsitic generation process to a stochastic one by injecting noise at each step of the flow integration. The update rule is defined as:
$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\left[{\bm{v}}_{\bm{\theta}}(\mathbf{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{x}_{t}+(1-t){\bm{v}}_{\theta}(\mathbf{x}_{t},t)\right)\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\bm{\epsilon},$ (4)
@@
Reward
To further refine the model’s output, we propose an online reinforcement learning (RL) phase subsequent to the initial RF training. This stage employs a bespoke policy-gradient algorithm to fine-tune the generative model $\mathbf{v}_{\bm{\theta}}$. The objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of video quality and identity consistency metrics:
$\bm{\theta}^{*}=\arg\max_{\bm{\theta}}\mathbb{E}_{\mathbf{V}\sim p_{\bm{\theta}}}\left[\mathcal{R}_{\text{total}}(\mathbf{V})\right],$ (5)
where $\mathcal{R}_{\text{total}}(\mathbf{V}):=\lambda_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V})+\lambda_{\text{id}}\mathcal{R}_{\text{id}}(\mathbf{V},\mathcal{I})$. Here, $\mathcal{R}_{\text{qual}}$ assesses the perceptual quality of the video, while $\mathcal{R}_{\text{id}}$ quantifies the identity preservation with respect to the reference images $\mathcal{I}$. The weights $\lambda_{\text{qual}}$ and $\lambda_{\text{id}}$ balance these two objectives.
@@ 3.4 Training Data curation
Figure 3: Data curation pipeline of ID-Composer.
Current subject-to-video generation methods are hampered by notable deficiencies in output quality and diversity. These shortcomings are largely attributable to the restrictive nature of paired training data, which proves inadequate for modeling complex, real-world variations in subject motion, camera viewpoint, and scene layout. To surmount these limitations, we propose a sophisticated data-curation pipeline that harnesses the capabilities of modern Vision-Language Models (VLMs) (Bai et al., 2025) and powerful image editing models (Fortin et al., 2025). As shown in Fig. 3, the culmination of this pipeline is a large-scale, composite dataset meticulously constructed to propel the development of next-generation multi-subject video synthesis models. This dataset is composed of three heterogeneous data sources. The first component consists of subject-video pairs extracted from a diverse collection of OpenS2V-Nexus (Yuan et al., 2025), offering a wide array of authentic scenes and actions. The second component comprises synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, systematically increasing the diversity of subject-background compositions. The third component is a high-fidelity collection of professionally shot videos with detailed annotations. More technical details are provided in Appendix C.
@@ 4.1 Experimental Settings
Our model, ID-Composer, is initialized from the weights of the Wan-Video-1.3 B model (Wan, 2025). To leverage superior semantic understanding, we employ a dual text-encoder architecture, combining T5 (Raffel et al., 2020) with Qwen2.5-VL-7B-Instruct (Bai et al., 2025). The model was trained for 30,000 iterations on our curated dataset at a resolution of 480p, utilizing 16 H20 GPUs. For inference, we employ an Euler sampler with 50 steps and classifier-free guidance (Ho & Salimans, 2022) to modulate the influence of image and text conditions, setting the classifier-free guidance scale to 2.5.
@@
We benchmark our method against state-of-the-art open-source models that support native subject-to-video generation. This includes Phantom (Liu et al., 2025b) and SkyReels-A2 (Fei et al., 2025), both available in 1.3B and 14B parameter variants, and VACE (Jiang et al., 2025). For a comprehensive assessment, we also present a qualitative comparison with leading proprietary systems, such as VIDU (Vidu, 2024), Pika (Pika, 2024) and Kling (keling, 2024).
@@
Following the benchmark protocol of OpenS2V-Nexus (Yuan et al., 2025), we perform evaluation on a diverse dataset of 180 unique subject-text pairs. We report on a suite of automated metrics, including Aesthetics, Motion quality, Face Similarity (FaceSim), and alignment scores against three benchmarks (GmeScore, NexusScore, NaturalScore). The Total Score is a normalized weighted sum of these metrics. Further details are provided in Appendix D.
@@ 4.2 Main Results
As demonstrated in Table 1, ID-Composer achieves state-of-the-art performance across all evaluation metrics, outperforming existing open-source models by a significant margin. Notably, our method shows substantial improvements in FaceSim and NexusScore, which we attribute to our hierarchical identity-preserving attention mechanism and the rich semantic guidance from the VLM encoder. These components work in concert to ensure high-fidelity identity preservation and robust alignment with complex textual prompts.
Table 1: Quantitative Comparison against existing methods for the open-domain subject-to-video benchmark. Total score is the normalized weighted sum of other scores. $\uparrow$” higher is better.
[…]
Proprietary Models
Vidu 2.0 (Vidu, 2024) 47.59% 41.47% 13.52% 35.11% 67.57% 43.55% 71.44%
Pika 2.1 (Pika, 2024) 48.88% 46.87% 24.70% 30.80% 69.21% 45.41% 69.79%
Kling 1.6 (keling, 2024) 54.46% 44.60% 41.60% 40.10% 66.20% 45.92% 79.06%
Open-source Models
VACE-1.3B (Jiang et al., 2025) 45.53% 48.24% 18.83% 20.58% 71.26% 37.95% 71.78%
VACE-14B (Jiang et al., 2025) 52.87% 47.21% 15.02% 55.09% 67.27% 44.20% 72.78%
Phantom-1.3B (Liu et al., 2025b) 50.71% 46.67% 14.29% 48.55% 69.43% 42.44% 70.26%
Phantom-14B (Liu et al., 2025b) 52.32% 46.39% 33.42% 51.48% 70.65% 37.43% 68.66%
SkyReels-A2-P14B (Fei et al., 2025) 49.61% 39.40% 25.60% 45.95% 64.54% 43.77% 67.22%
Ours-1.3B (Base Model) 54.33% 42.50% 38.00% 58.12% 64.00% 43.00% 71.00%
Ours-14B 57.05% 45.00% 40.00% 60.50% 67.00% 45.00% 73.00%
@@
Figure 4 presents a qualitative comparison against leading methods. ID-Composer consistently generates videos with superior visual quality, temporal coherence, and identity preservation. While baseline methods often exhibit artifacts or identity drift, especially in longer videos or complex scenes, our model maintains a stable representation of all subjects , accurately reflecting their appearance from the reference images.
Figure 4: Qualitative Comparison with State-of-the-Art Methods. Our method, ID-Composer, demonstrates superior performance in identity preservation, temporal consistency, and alignment with the textual prompt compared to existing open-source and proprietary models.
@@ 4.3 Ablation and Analysis
We conduct a series of ablation studies to validate the effectiveness of our key design choices. The results, summarized in Table 2, underscore the importance of each proposed component.
Table 2: Ablation studies of ID-Composer. We evaluate the impact of our key contributions: the hierarchical identity-preserving attention, the VLM for semantic understanding, and our curated dataset. All metrics are reported on our evaluation set. $\uparrow$” indicates that higher is better.
[…]
ID-Composer (Full Model) 58.12% 49.55% 48.91% 54.33%
w/o Hierarchical Attention 51.34% 48.98% 47.52% 50.11%
w/o VLM Encoder 56.98% 42.17% 46.88% 49.89%
w/o Curated Data 54.55% 45.32% 45.13% 48.78%
@@
We evaluated the impact of the VLM by replacing the dual-encoder setup with a single T5 encoder (Raffel et al., 2020). The results (“w/o VLM Encoder”) show a marked decrease in Text-Video Alignment. This highlights the VLM’s superior ability to parse complex, multi-subject prompts and provide fine-grained semantic guidance, which is essential for accurate video generation.
@@
Finally, we trained the full model on a baseline dataset without our curated data. This version (”w/o Curated Data”) yielded lower scores across all metrics, particularly in Video Quality. This demonstrates that our data curation pipeline, which synthesizes diverse and high-quality training examples, is vital for achieving robust and generalizable performance. Qualitative results in Figure 5 further illustrate these findings.
@@
Figure 5: Qualitative Ablation Study. The left panel highlights the importance of our curated dataset, showing improved coherence and realism in subject integration compared to a model trained without it. The right panel illustrates the effectiveness of our online reinforcement learning stage, which significantly enhances visual quality and subject consistency.
4.4 Application: Controllable Video Editing
The architecture of ID-Composer naturally lends itself to a range of editing applications. By providing an existing video as an initial state and conditioning the generation on new subjects or modified text prompts, our model can perform zero-shot, identity-consistent video editing. As illustrated in Figure 6, ID-Composer can seamlessly insert new subjects into a scene, replace existing ones, or alter the background, all while maintaining temporal and semantic coherence. This capability opens up new avenues for personalized content creation and virtual storytelling.
@@ 4.4 Application: Controllable Video Editing
Figure 6: Applications in Controllable Video Editing. ID-Composer enables zero-shot editing of existing videos, including subject replacement and background modification, while preserving identity and temporal consistency.

### v3

@@ 3.3 Post-training with Online RL
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim P(Q),\\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\end{subarray}}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}$ (3)
$\quad\left\{\min\left[r_{t}^{i}(\theta)\hat{A}_{i,t},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]\right$.
$\quad\left.-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\,\|\,\pi_{\text{ref}}\big]\right\},$
$\text{where}\quad r_{t}^{i}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
@@
the relative quality $\hat{A}_{i,t}$ of the i-th response is computed as $\hat{A}_{i,t}=\frac{r_{i}-\text{Mean}(\{r_{1},r_{2},\ldots,r_{n}\})}{\text{Std}(\{r_{1},r_{2},\ldots,r_{n}\})},$
@@
While convert flow matching’s determinsitic generation process to a stochastic one by injecting noise at each step of the flow integration. The update rule is defined as:
$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}$ $+\left[{\bm{v}}_{\bm{\theta}}(\mathbf{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{x}_{t}+(1-t){\bm{v}}_{\theta}(\mathbf{x}_{t},t)\right)\right]\Delta t$ (4)
$+\sigma_{t}\sqrt{\Delta t}\,\bm{\epsilon}$
@@
Reward Design.
The objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of video quality and identity consistency:
$\mathcal{R}_{\text{total}}(\mathbf{V}):=w_{\text{fid}}\mathcal{R}_{\text{fid}}(\mathbf{V},\mathcal{I})+w_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V}),$ (5)
where we empirically set the weights to $w_{\text{fid}}=0.6$, $w_{\text{qual}}=0.4$ to prioritize identity preservation while ensuring high quality and text alignment.
Fidelity Reward ($\mathcal{R}_{\text{fid}}$).
The fidelity reward ensures that the generated subjects are consistent with the reference images $\mathcal{I}$. It combines a face-specific metric with a holistic subject consistency score:
$\mathcal{R}_{\text{fid}}:=(1-\alpha)\mathcal{R}_{\text{face}}+\alpha\mathcal{R}_{\text{subject}}.$ (6)
For facial identity $\mathcal{R}_{\text{face}}$, We use a robust combination of the mean and minimum FaceSim scores [consisid] across all $N$ subjects:
$\mathcal{R}_{\text{face}}:=(1-\gamma)\left(\frac{1}{N}\sum_{k=1}^{N}\mathcal{R}_{\text{id}}^{k}\right)+\gamma\left(\min_{k=1,\dots,N}\mathcal{R}_{\text{id}}^{k}\right),$ (7)
where $\mathcal{R}_{\text{id}}^{k}$ is the ArcFace-based score for subject $k$, and we set $\gamma=0.5$. For overall subject consistency, $\mathcal{R}_{\text{subject}}$, we employ the NexusScore, a metric from our proposed OpenS2V-Eval benchmark that assesses non-facial and holistic subject attributes. We balance face and subject with $\alpha=0.5$.
Quality Reward ($\mathcal{R}_{\text{qual}}$).
The quality reward assesses both the aesthetic appeal and the physical plausibility of the generated video $\mathbf{V}$:
$\mathcal{R}_{\text{qual}}:=(1-\beta)\mathcal{R}_{\text{aes}}+\beta\mathcal{R}_{\text{nat}}.$ (8)
$\mathcal{R}_{\text{aes}}$ is the standard Aesthetic Score [improved-aesthetic-predictor]. To combat reward hacking that may produce unnatural but aesthetically pleasing artifacts, we introduce $\mathcal{R}_{\text{nat}}$. This is the NaturalScore from our benchmark, which leverages a powerful VLM to evaluate whether the video adheres to physical laws and common sense. We set $\beta=0.4$ to strongly penalize unnatural generations.
Figure 3: Data curation pipeline of ID-Crafter.
@@ 3.4 Training Data curation
Current subject-to-video generation methods are hampered by notable deficiencies in output quality and diversity. These shortcomings are largely attributable to the restrictive nature of paired training data, which proves inadequate for modeling complex, real-world variations in subject motion, camera viewpoint, and scene layout. To surmount these limitations, we propose a sophisticated data-curation pipeline that harnesses the capabilities of modern VLMs (e.g., QwenVL-72B) [Qwen2.5-VL] and a powerful image editing model (a.k.a., Nano Banana) [nanobanana]. The first component consists of subject-video pairs extracted from a diverse collection of OpenS2V-Nexus [yuan2025opens2v], offering a wide array of authentic scenes and actions. The second component comprises synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, including explicitly designed cross-subject compositions and fusion examples, to mitigate “copy-plate” artifacts and enhance multi-entity interaction coherence. The third component is a high-fidelity collection of professionally shot videos with detailed annotations. As shown in Fig. 3, the culmination of this pipeline is a large-scale, composite dataset meticulously constructed to propel the development of next-generation multi-subject video synthesis models. This dataset is composed of three heterogeneous data sources. More technical details are provided in Appendix.
@@ 4.1 Experimental Settings
[…]
Proprietary Models
Vidu 2.0 [Vidu] 47.59% 41.47% 13.52% 35.11% 67.57% 43.55% 71.44%
Pika 2.1 [Pika] 48.88% 46.87% 24.70% 30.80% 69.21% 45.41% 69.79%
Kling 1.6 [Keling] 54.46% 44.60% 41.60% 40.10% 66.20% 45.92% 79.06%
Open-source Models
VACE-1.3B [VACE] 45.53% 48.24% 18.83% 20.58% 71.26% 37.95% 71.78%
VACE-14B [VACE] 52.87% 47.21% 15.02% 55.09% 73.27% 44.20% 72.78%
Phantom-1.3B [liu2025phantom] 50.71% 46.67% 14.29% 48.55% 69.43% 42.44% 70.26%
Phantom-14B [liu2025phantom] 52.32% 46.39% 33.42% 51.48% 70.65% 37.43% 68.66%
SkyReels-A2-P14B [fei2025skyreels] 49.61% 39.40% 25.60% 45.95% 64.54% 43.77% 67.22%
Ours-1.3B (Base Model)
$\star$
54.33% 42.50% 38.00% 58.12% 64.06 % 43.22% 71.09%
Ours-1.3B (Base Model + online RL)
$\star$
55.16% 48.85% 36.50% 66.10% 62.20% 43.45% 69.15%
Ours-14B (Large-scale Model) 57.05% 45.28% 40.34% 60.71% 67.49% 45.11% 73.23%
Table 1: Quantitative Comparison against existing methods for the open-domain subject-to-video benchmark. Total score is the normalized weighted sum of other scores. $\uparrow$” indicates that higher is better.
@@
Our model, ID-Crafter, is initialized from the weights of the Wan-Video-1.3 B model [wan2025]. To leverage superior semantic understanding, we employ a dual text-encoder architecture, combining T5 [T5] with Qwen2.5-VL-7B-Instruct [Qwen2.5-VL]. The model was trained for 30,000 iterations on our curated dataset at a resolution of 480p, utilizing 16 H20 GPUs. For inference, we employ an Euler sampler with 50 steps and classifier-free guidance [ho2022classifier] to modulate the influence of image and text conditions, setting the classifier-free guidance scale to 2.5. The inference time for our 1.3B model is approximately one minute to generate a video at 480p resolution.
@@
We benchmark our method against state-of-the-art open-source models that support native subject-to-video generation. This includes Phantom [liu2025phantom] and SkyReels-A2 [fei2025skyreels], both available in 1.3B and 14B parameter variants, and VACE [VACE]. For a comprehensive assessment, we also present a qualitative comparison with leading proprietary systems, such as VIDU [Vidu], Pika [Pika] and Kling [Keling].
@@
Following the benchmark protocol of OpenS2V-Nexus [yuan2025opens2v], we perform evaluation on a diverse dataset of 180 unique subject-text pairs. To ensure a rigorous and unbiased evaluation, this dataset is maintained as a held-out set, completely disjoint from our training data, thereby preventing any potential data leakage. We report on a suite of automated metrics, including Aesthetics, Motion quality, Face Similarity (FaceSim), and alignment scores against three benchmarks (GmeScore, NexusScore, NaturalScore). In our ablation studies, we also report the VLM-based visual scorer, Q-Align [wu2023qalign], for perceptual quality.
@@ 4.2 Main Results
As demonstrated in Table 1, ID-Crafter achieves state-of-the-art performance across all evaluation metrics, outperforming existing open-source models by a significant margin. Notably, our method shows substantial improvements in FaceSim and NexusScore, which we attribute to our hierarchical identity-preserving attention mechanism and the rich semantic guidance from the VLM encoder. These components work in concert to ensure high-fidelity identity preservation and robust alignment with complex textual prompts.
@@
Fig. 4 presents a qualitative comparison against leading methods. ID-Crafter consistently generates videos with superior visual quality, temporal coherence, and identity preservation. While baseline methods often exhibit artifacts or identity drift, especially in longer videos or complex scenes, our model maintains a stable representation of all subjects , accurately reflecting their appearance from the reference images.
Figure 4: Qualitative Comparison with State-of-the-Art Methods. Our method, ID-Crafter, demonstrates superior performance in identity preservation, temporal consistency, and alignment with the textual prompt compared to existing open-source and proprietary models.
@@ 4.3 Ablation and Analysis
We evaluated the impact of the VLM by replacing the dual-encoder setup with a single T5 encoder [T5]. The results (“w/o VLM Encoder”) show a marked decrease in Text-Video Alignment. This highlights the VLM’s superior ability to parse complex, multi-subject prompts and provide fine-grained semantic guidance, which is essential for accurate video generation.
@@
Finally, we trained the full model on a baseline dataset without our curated data. This version (“w/o Curated Data”) yielded lower scores across all metrics, particularly in Video Quality. Critically, the base model exhibited more severe “copy-plate” artifacts, due to limited exposure to cross-subject interactions. Our data curation pipeline explicitly mitigates this by incorporating synthetic cross-subject compositions and fusion examples, which enhances the model’s ability to synthesize coherent multi-entity interactions. This demonstrates that our data curation pipeline, which synthesizes diverse and high-quality training examples, is vital for achieving robust and generalizable performance. Qualitative results in Fig. 5 further illustrate these findings.
Table 2: Ablation studies on the key components of ID-Crafter. We compare the full model against variants removing hierarchical attention, the VLM encoder, and the curated dataset.
Method FaceSim $\uparrow$ Q-Align $\uparrow$ Video Quality $\uparrow$ Total Score $\uparrow$
[…]
$\star$
58.12% 0.351 48.91% 54.33%
   w/o Hierarchical Attention 51.34% ($\downarrow$11.7%) 0.348 ($\downarrow$0.9%) 47.52% ($\downarrow$2.8%) 50.11% ($\downarrow$7.8%)
   w/o VLM Encoder 56.98% ($\downarrow$2.0%) 0.287 ($\downarrow$18.2%) 46.88% ($\downarrow$4.2%) 49.89% ($\downarrow$8.2%)
   w/o Curated Data 54.55% ($\downarrow$6.1%) 0.321 ($\downarrow$8.5%) 45.13% ($\downarrow$7.7%) 48.78% ($\downarrow$10.2%)
@@
Figure 5: Qualitative Ablation Study. The left panel highlights the importance of our curated dataset, demonstrating improved coherence and realism in subject integration (e.g., mitigating ‘copy-paste’ artifacts) compared to a model trained without it. Meanwhile, the right panel illustrates the effectiveness of our online reinforcement learning stage, which significantly enhances visual quality and subject consistency.
@@
Figure 6: Applications in Controllable Video Editing. ID-Crafter enables zero-shot editing of existing videos, including subject replacement and background modification, while preserving identity and temporal consistency.
Effect of Online RL
As shown in Tab. 3, the online RL stage brings substantial improvements across all metrics compared to the SFT baseline. Specifically, our model achieves a 13.7% relative increase in FaceSim and a 14.9% relative increase in Aesthetics. The alignment score, Q-Align, also sees a significant 16.8% relative improvement. This demonstrates that directly optimizing the model with our composite reward function is crucial for enhancing identity preservation and perceptual quality. The qualitative results in Fig. 5 (right panel) further corroborate this finding, where the model trained with RL generates videos with significantly better subject consistency compared to the SFT baseline.
Table 3: Analysis of the online RL optimization. We compare our online GRPO approach with offline DPO and a standard SFT baseline. We also ablate the components of our composite reward function $\mathcal{R}_{\text{total}}$.
Method FaceSim $\uparrow$ Aesthetics $\uparrow$ Q-Align. $\uparrow$ Total Score $\uparrow$
SFT Baseline 58.12% 42.50% 0.351 54.33%
Comparison of RL Algorithms
DPO (Offline) 62.35% 45.15% 0.382 54.80%
Ours (Online) 66.10% 48.85% 0.410 55.16%
Ablation on Reward Components
Ours w/o Fidelity ($\mathcal{R}_{\text{fid}}$) 45.32% 47.10% 0.391 53.50%
Ours w/o Quality ($\mathcal{R}_{\text{qual}}$) 63.50% 43.81% 0.379 54.82%
Ours w/o Natural ($\mathcal{R}_{\text{nat}}$) 69.30% 50.83% 0.361 53.01%
Ours (Online) 66.10% 48.85% 0.410 55.16%
Analysis of Online RL Optimization
Offline RL methods like DPO [DPO] offer an alternative but rely on a static dataset of preference pairs, which may not sufficiently cover the vast space of possible generations. In contrast, our online approach allows the policy to actively explore and receive feedback from our reward model, leading to more effective optimization. As shown in Tab. LABEL:↑:rl_ablation, our online GRPO method significantly outperforms both the SFT baseline and the offline approach across all metrics. This confirms that online sample generation is more effective for fine-tuning video generation models. The table also presents an ablation study on the components of our composite reward function, $\mathcal{R}_{\text{total}}$. Removing the fidelity term ($\mathcal{R}_{\text{fid}}$) causes a severe drop in FaceSim, while removing the quality ($\mathcal{R}_{\text{qual}}$) term leads to a degradation in Aesthetics. Interestingly, removing the Natural Reward $\mathcal{R}_{\text{nat}}$ leads to a degradation in Q-Align, but a slight increase in FaceSim and Aesthetics. This suggests a reward hacking scenario, where the model over-optimizes for certain metrics at the expense of overall alignment. These results validate our multi-faceted reward design and the necessity of each component for achieving a balanced improvement.
Controllable Video Editing
The architecture of ID-Crafter naturally lends itself to a range of editing applications. By providing an existing video as an initial state and conditioning the generation on new subjects or modified text prompts, our model can perform zero-shot, identity-consistent video editing. As illustrated in Fig. 6, ID-Crafter can seamlessly insert new subjects into a scene, replace existing ones, or alter the background, all while maintaining temporal and semantic coherence. This capability opens up new avenues for personalized content creation and virtual storytelling.